Circular No. 5795

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

68 OPHIUCHI A. Vidal-Madjar, R. Ferlet, A. Lecavelier des Etangs, G. Perrin, and F.

Sèvre, Institut d’Astrophysique de Paris; F. Colas, Bureau des Longitudes; F. Deladerrière, Observatoire du Pic du Midi; and A.-M. Lagrange-Henri, Observatoire de Grenoble, communicate: “Observations with the Bernard Lyot 2-m telescope at the Pic du Midi Observatory on 1992 Aug. 26 revealed a disk-like structure around the nearby main-sequence star 68 Oph (HR 6723, type A2 Vn, distance 48 pc, m_{v} = 4.45). The observations were done in the 695-nm band with an ‘anti-blooming’ CCD camera. The disk is viewed nearly edge-on between 1′′.5 and 7′′ from the star, the limit of our field-of-view. It thus extends out to at least 350 AU. The disk is inclined by 13^{o} to the east of north, and its magnitude ranges from 16 to 22 arcsec^{−2}. The radial luminosity of the disk is well fit by an r^{−3} power law out to 3′′ on both sides, and is separated then into an r^{−2} power law in the northern extension and an r^{−4} power law in the southern one. 68 Oph is known as a member of a double system whose companion is a 6.9-magnitude star with a separation of 0′′.6 in p.a. 70^{o}. In addition to the strong asymmetry observed in the disk, a slight wrapping is also detected at more than 5′′ from the star — possible signatures of the companion interaction. Following β Pic (Smith and Terrile 1984, Science 226, 1421), this is the second disk to be directly imaged optically, and it is the first one related to a double system. Due to the high S/N level (30σ at 3′′ from the star) and to the remarkable similarity to the β Pic disk, the detection seems to be quite convincing, although 68 Oph does not present any infrared excess in the IRAS Point Source Catalogue. We strongly recommend other observations to confirm this detection.”

SUPERNOVA 1993J IN NGC 3031 M. Kidger, Instituto de Astrofisica de Canarias, reports the following

infrared photometry obtained by T. Mahoney, M. Selby, and himself using the 4.2-m William Herschel Telescope at Roque de los Muchachos Observatory (6′′.7 aperture, approximately east-west 10′′ chop; calibration star BS 3403, previously calibrated against Vega): Apr. 29.818 UT, K = 10.01 ± 0.04; May 1.837, K = 10.20 ± 0.03; 5.840, K = 10.16 ± 0.04, H = 10.32 ± 0.04; 6.833, K = 10.35 ± 0.04, H = 10.53 ± 0.04, J = 10.87 ± 0.04. The quoted errors are provisional, pending reanalysis.

1993 May 18 Daniel W. E. Green